1. Discrete distribution for X is given by the following table: Probabilities p Values A 0.2 20 0.2 40 0.3 60 0.2 80 0.1 100 Find distribution function fax) and median Me(X). Calculate expectation value MX), variance (dispersion) DA), standard error σ(X), asymmetry coefficient As(X) and excess Ex(X). 2. Calculate multiplier k. Find distribution function foc), mode Moc), median Me(x), expectation value M(x), variance (dispersion) D(x), standard error σ(x), asymmetry coefficient As(x) and excess Ex(x) for continuous distributions with the given probability densities a) b) 0 r<-2 -2 r>6 0sxs8 ρ(x) = k(4x-x2 +12) 6 x 1r+1 8s10 10 x>10 Calculate probability that X E 1-28 Calculate probability that x e -3:2]

This is a statistics question, unfortunately I have difficulties understanding problems with statistics and I need this question solved with full workings and also in an understandable way to revise. Your help will be highly appreciated, *I will give positive rating to the fullest*. Thank you!!
Answer #1

1) f(X)

X f(X= x)
20 0.2
40 0.2
60 0.3
80 0.2
100 0.1
1

Median

P(X <median) = 0.5

X f(X= x) cumulative probability
20 0.2 0.2
40 0.2 0.4
60 0.3 0.7
80 0.2 0.9
100 0.1 1

note that  

P(X< = 40) = 0.4 and P(X<= 60) = 0.7

as 0.5 is between 0.4 and 0.7

60 is median

M(X)=Xp(X=z)X = 20 * 0.2+40 * 0.2 +60 * 0.3480 * 0.2 + 100 * 0.1

= 56

D(X)= E(X^2) - (E(X))^2

X f(X= x) x*p x^2*p
20 0.2 4 80
40 0.2 8 320
60 0.3 18 1080
80 0.2 16 1280
100 0.1 10 1000
1 56 3760

E(X^2) = 3760

hence

D(X) = 3760 - 56^2

= 624

standard error = sd(X) = sqrt(D(X)) = sqrt(624) = 24.979999

he most frequently employed measure of the asymmetry of a distribution, defined by the relationship

11 = S/2

where a0135902.png and a0135903.png are the second and third central moments of the distribution, respectively

X f(X= x) x*p x^2*p (x - mu)^3
20 0.2 4 80 -46656
40 0.2 8 320 -4096
60 0.3 18 1080 64
80 0.2 16 1280 13824
100 0.1 10 1000 85184
1 56 3760 48320

As(X) =   48320/(3760)^(3/2)

= 0.2095777

A scalar characteristic of the pointedness of the graph of the probability density of a unimodal distribution. It is used as a certain measure of the deviation of the distribution in question from the normal one. The excess e0368001.png is defined by the formula

e0368002.png

where e0368003.png is the second Pearson coefficient (cf. Pearson distribution), and e0368004.png and e0368005.png are the second and fourth central moments of the probability distribution

X f(X= x) x*p x^2*p (x - mu)^3 (x- mu)^4
20 0.2 4 80 -46656 1679616
40 0.2 8 320 -4096 65536
60 0.3 18 1080 64 256
80 0.2 16 1280 13824 331776
100 0.1 10 1000 85184 3748096
1 56 3760 48320 5825280

Ex(X) = 5825280/3760^2 - 3

= -2.58796